(1 pt) Consider a list of randomly generated 3-letter "words" printed on a paper. The letters cannot be repeated.
(a) At least how many of these "words" should be printed to be sure of having at least 8 identical "words" on the list?
Answer =
(b) At least how many identical "words" are printed if there are 140401 "words" on the list?
Answer =

Answer by sudhanshu_kmr(1152)

no. of possible 3-letter worlds = 26*25*24 =15600

(a) no. of minimum words to sure having 8 identical words = 15600*7 +1 =109201


(b) divide 140401 by 15600
we get 140401 = 15600 * 9 + 1
so, at least 10 identical words.



It is possible that some typing mistake in solution of a problem, if any please ignore it. Understand the concept and try to solve the problem yourself. If there is problem related to concept, contact at
sudhanshu.cochin@yahoo.com or sudhanshu.cochin@gmail.com
Best of luck.......